a balloon takes up 625L of helium at STP, if it is heated to 80.1°C, what will its new volume be to make it pressure constant
1
Expert's answer
2020-03-27T09:08:48-0400
"\frac{V_1}{T_1}=\frac{V_2}{T_2}"

"\frac{625}{273.15}=\frac{V_2}{80.1+273.15}"

"V_2=808\ L"